OpenStudy (billyjean):

How is a projectile’s initial speed related to its height and horizontal displacement?

OpenStudy (billyjean):

@paki

Parth (parthkohli):

Well, what do you know about projectiles to begin with?

OpenStudy (billyjean):

I only know that a projectile is a missile designed to be fired from a rocket or gun.

Parth (parthkohli):

So the math part of it?

OpenStudy (billyjean):

i dont know im confused

Parth (parthkohli):

Well, alright. So do you know equations of motion?

OpenStudy (billyjean):

no

Parth (parthkohli):

Aw man, I can't help you out then

OpenStudy (billyjean):

why

Parth (parthkohli):

you need to be familiar with the basics to figure this out

OpenStudy (billyjean):

v = u + at From acceleration definition s = ½ (v+u) × t Average velocity is ½ (v+u) s = ut + ½ at² No 'v' needed v² = u² + 2as No 't' needed

OpenStudy (billyjean):

is that it

OpenStudy (billyjean):

@3mar

OpenStudy (3mar):

Oh I was checking that question before and have prepared something for it, but it was closed!

OpenStudy (bugman954):

Simple the greater the speed the greater the height and horizontal displacement

OpenStudy (vipulrakta):

 and the relation between u; h , r is\[H = (u ^{2}\sin ^{2} \theta) / 2g , R=(u^{2} \sin 2\theta )/g\] grater the initial speed more is horizantal distance and hight is more ance vice versa . R is maximum when angle = 45